**Please use numbers given** Please answer the blank boxes in (a) and (b) - thank you!! :)

Standard Product Cost, Direct Materials Variance Condiments Company uses standards to control its materials costs. Assume thaa. Determine the standard unit materials cost per pound for a standard batch. If required, round amounts to the nearest cent.
Answer #1

a) Calculate standard cost per batch

Standard cost per batch
Whole tomatoes 468
Vinegar 154
Corn syrup 48
Salt 56
Total 726
Standard unit material cost per pound (726/800) 0.91

b) Material quantity variance

Material quantity variance Favorable/Unfavorable
Whole tomatoes (1300-1400)*0.36 = 36 Unfavorable
Vinegar (70-67)*2.2 = 6.60 Favorable
Corn syrup (6-7)*8 = 8 Unfavorable
Salt (28-27)*2 = 2 Favorable
Total direct material quantity variance 35.40 Unfavorable